An unstretched spring is 10 ft long. A pull of 40 lb stretches the

spring by ½ ft. Find the work done in stretching the spring from

10 ft to 14 ft.


1
Expert's answer
2022-04-27T11:49:10-0400

Hooke's Law


"F=-k x"

"k=\dfrac{40\ lb}{1/2\ ft}=80\ lb/ft"

The work done in stretching the spring is


"W=k(x_2-x_1)^2\/2"

"W=80\ lb/ft\cdot(14\ ft-10\ ft)^2/2=640\ ft- lb"